OpenStudy (anonymous):

Find the product of (x + 7)2. x2 − 14x + 49 x2 + 14x + 49 x2 − 49 x2 + 49 i got tthe scond one am i right?

OpenStudy (saifoo.khan):

expand (x+7)(x+7)

OpenStudy (lgbasallote):

(a+b)^2 = a^2 + 2ab + b^2 where your a = x; b = 7...just substitute

OpenStudy (anonymous):

so x2+14+49?

OpenStudy (anonymous):

its \[x ^{2}+14x+49\] you're correct